The noise budget is updated and plotted with today's beat measurement.
(*The electronic noise plotted in the graph is not correctly calibrated, see psl:816 for the complete calculation)
After we replaced the table's broken leg, we floated the table and measured the beat signal as a reference before modifying the seismic stack. The calculation agree with the measured data quite well.
We also measured the electronic noise from PLL. This was the signal which was fed back to the LO (with SR560 gain = 20). Apparently, we are sitting on it at 1kHz and above. We definitely need to work on PLL readout system to measure at lower sensitivity.
New traces in the noise budget:
I removed noises from room temperature fluctuation/ heater noise and spacer thermal noise because they are way smaller than coating noise and crowd the plot.
==Details about some traces in the noise budget==
The vertical seismic noise coupling is calculated by applying the seismic measurement times stack transfer function times cavity bending coupling
[ Frequency noise from seismic ] = [measured data] x [stacks TF] x [bending factor].
The peak around 6 Hz might be coupled from horizontal direction. This will be added soon.
RIN induced length noise is still an estimated. We have not been able to measured the real coupling yet, as the SNR is so low.
LO phase noise: This is from measurement. I'm not quite sure if I miss some calibration factors. The phase noise does not show up in the beat yet even though it is very close together right now.